Suppose that a random sample of size n=100 is selected from a binomial population proportion with the population proportion 0.25
The calculated standard error equal to
1. 0.0019
2. 0.0433
3. 0.0025
4. 0.0001
5. 25
1
Expert's answer
2020-10-29T20:05:21-0400
"\mu_{\hat{p}}=p=0.25"

"\sigma_{\hat{p}}=\sqrt{\dfrac{p(1-p)}{n}}=\sqrt{\dfrac{0.25(1-0.25)}{100}}\approx 0.0433"

The answer is 2. 0.0433.